OpenStudy (anonymous):

A high school chess club has 12 members. In how many different ways can they choose a president and vice president? A. 144 B. 66 C. 23 D. 132

OpenStudy (solomonzelman):

In this case it would be permutations, because you when you choose person D and person E they can be rearranged in 2 different ways, D is ores and E is vice or the other way. \(\Large\color{blue}{ ~_{12}~P~~_2~=~ \frac{12!}{(12-2)!} }\)

OpenStudy (anonymous):

@SolomonZelman can i just have the answer please

OpenStudy (solomonzelman):

\(\Large\color{black}{ \frac{12!}{(12-2)!} =\frac{12!}{10!}=\frac{10!\times 11 \times 12}{10!} =11 \times 12=?}\)

OpenStudy (solomonzelman):

just do the last step:)